# Chao-Han Huck Yang1, Ramani Duraiswami2, Dinesh Manocha2, Rafael Valle1, Bryan Catanzaro1

### v1

@@ A.2 Training Datasets
GigaSpeech (L) (chen2021gigaspeech, ) 2,266,371
Speech-in-Sound Captions abu2016youtube 1,999,959
SPGISpeech (o2021spgispeech, ) 1,966,109
Sound-VECaps yuan2025sound 1,657,029
Million Songs Dataset bertin2011million 1,169,997
Common Voice 15 commonvoice:2020 1,109,689
MiraData ju2024miradata 748,320
Action2sound chen2024action2sound 306,602
NSynth engel2017neural 289,205
LibriSpeech (panayotov2015librispeech, ) 281,241
Freesound fonseca2017freesound 256,695
AudioSet Strong hershey2021benefit 216,622
VGGSound chen2020vggsound 185,161
VoxPopuli (en) (wang2021voxpopuli, ) 177,019
FMA defferrard2016fma 106,412
Video Recap islam2024video 64,627
CochlScene jeong2022cochlscene 60,855
Music4All (santana2020music4all, ) 109269
Switchboard (godfrey1992switchboard, ) 76,652
FSD50k fonseca2021fsd50k 40,966
MACS morato2021diversity 31,675
@@
MagnaTagATune law2010evaluation 25,863
SoundDescs koepke2022audio 23,085
Clotho drossos2020clotho 19,195
TAU-Urban mesaros2018multi 14,400
MusicCaps agostinelli2023musiclm 5,479
WavText5K deshmukh2022audio 4,347
SONICS rahman2024sonics 1,602
@@
MUSDB18 rafii2017musdb18 276
Medleydb-Pitch bittner2014medleydb 103
@@ Appendix B AudioSkills-XL
Temporal 188K Table 14 in (kong2025audioflamingo2, ) Table 14 in (kong2025audioflamingo2, ) Table 14 in (kong2025audioflamingo2, )
@@
Attribute Identification 201K Table 14 in (kong2025audioflamingo2, ) Table 14 in (kong2025audioflamingo2, ) Table 14 in (kong2025audioflamingo2, )
Counting 50K Table 14 in (kong2025audioflamingo2, ) Table 14 in (kong2025audioflamingo2, ) Table 14 in (kong2025audioflamingo2, )
Contextual Sound Event Reasoning 982K Table 14 in (kong2025audioflamingo2, ) Table 14 in (kong2025audioflamingo2, ) Table 14 in (kong2025audioflamingo2, )
Contextual Speech Event Reasoning 1,272K Table 14 in (kong2025audioflamingo2, ) Table 14 in (kong2025audioflamingo2, ) Table 14 in (kong2025audioflamingo2, )
Information Extraction 858K Table 14 in (kong2025audioflamingo2, ) Table 14 in (kong2025audioflamingo2, ) Table 14 in (kong2025audioflamingo2, )
General Reasoning 704K Table 14 in (kong2025audioflamingo2, ) Table 14 in (kong2025audioflamingo2, ) Table 14 in (kong2025audioflamingo2, )
@@
Speech-in-Sound QA (ours) 1,739K Speech-in-Sound Caps (YouTube8M) Caption, Transcripts, Speech Characteristics Fig. 25 49
@@ E.2 Clustering for constructing AF-Chat
Specifically, for each seed audio, we retrieve its top 8 most semantically similar and top 8 most dissimilar clips from the dataset. Similarity is computed using captions, NV-Embed-v2 embeddings of the captions, and FAISS-based similarity search (douze2024faiss, ) of the embeddings.
@@
Order VoxPopuli (wang2021voxpopuli, ) 16,926 1.87 294.80 89.55
LibriSpeech panayotov2015librispeech 2,340 16.02 147.59 82.01
MELD poria2018meld 4,135 1.06 108.01 30.42
IEMOCAP (busso2008iemocap, ) 599 82.00 542.00 272.45
EuroParl (koehn2005europarl, ) 11,885 2.59 176.14 69.34
Fisher (cieri2004fisher, ) 25,962 33.34 240.00 136.84
Switchboard (godfrey1992switchboard, ) 2,702 22.81 148.96 87.38
MultiDialog (mesaros2018multi, ) 27,927 1.31 499.33 135.10
VoxCeleb2 (chung2018voxceleb2, ) 12,855 8.00 1273.60 71.12
Emotion Ident. IEMOCAP busso2008iemocap 300 82.00 542.00 272.22
MELD poria2018meld 1,847 1.78 108.01 33.20
Emotion Causal Reason. IEMOCAP busso2008iemocap 300 82.00 542.00 272.22
MELD poria2018meld 1,850 1.57 108.01 33.13
Emotion Flip Reason. IEMOCAP busso2008iemocap 299 82.00 542.00 272.62
MELD poria2018meld 1,807 1.64 108.01 33.57
Topic Relation. Reason. VoxPopuli wang2021voxpopuli 13,651 3.58 240.44 97.14
LibriSpeech panayotov2015librispeech 1,165 16.02 147.59 82.11
MELD poria2018meld 1,518 1.89 108.01 34.45
IEMOCAP busso2008iemocap 188 82.00 542.00 270.99
EuroParl koehn2005europarl 9,381 7.97 176.14 70.14
Fisher cieri2004fisher 20,453 33.34 240.00 136.10
Switchboard godfrey1992switchboard 998 24.58 148.96 90.05
MultiDialog mesaros2018multi 14,906 5.11 499.33 135.35
DailyTalk lee2023dailytalk 3,141 8.05 103.66 35.32
VoxCeleb2 chung2018voxceleb2 5,414 8.51 1193.60 78.96
Sarcasm Ident. IEMOCAP busso2008iemocap 299 82.00 542.00 271.58
MELD poria2018meld 1,958 1.10 108.01 31.82
Summarization VoxPopuli wang2021voxpopuli 13,913 2.12 294.80 91.38
LibriSpeech panayotov2015librispeech 1,057 16.02 147.59 83.15
MELD poria2018meld 2,803 1.84 108.01 32.92
IEMOCAP busso2008iemocap 300 82.00 542.00 272.22
EuroParl koehn2005europarl 8,905 6.62 176.14 70.03
Fisher cieri2004fisher 15,500 0.33 240.00 135.60
Switchboard godfrey1992switchboard 1,346 24.58 148.96 87.60
MultiDialog mesaros2018multi 20,838 1.93 499.33 135.73
DailyTalk lee2023dailytalk 7,218 8.05 103.66 31.42
VoxCeleb2 chung2018voxceleb2 5,894 7.94 1193.60 70.87
Spotify Podcasts clifton-etal-2020-100000 103920 0.06 18206.44 2002.99
Needle QA (IE) DailyTalk lee2023dailytalk 13,563 5.72 103.66 31.12
EuroParl koehn2005europarl 18,426 6.57 176.14 70.10
Fisher cieri2004fisher 37,779 18.59 240.00 135.99
IEMOCAP busso2008iemocap 542 82.00 542.00 272.03
LibriSpeech panayotov2015librispeech 2,248 16.02 147.59 82.82
Spotify Podcasts clifton-etal-2020-100000 103920 0.06 18206.44 2002.99
Response QA (IE) VoxPopuli wang2021voxpopuli 13,913 2.12 294.80 91.38
MELD poria2018meld 1,660 1.57 108.01 31.83
IEMOCAP busso2008iemocap 177 82.00 542.00 272.52
MultiDialog mesaros2018multi 13,505 1.95 499.33 135.40
DailyTalk lee2023dailytalk 4,516 5.72 103.66 30.91
Switchboard godfrey1992switchboard 862 22.81 148.96 88.75
@@
Causal QA (IE) VoxPopuli wang2021voxpopuli 12,264 4.10 240.44 92.88
LibriSpeech panayotov2015librispeech 1,166 16.02 147.59 82.04
MELD poria2018meld 2,957 1.27 108.01 31.74
IEMOCAP busso2008iemocap 298 82.00 542.00 273.10
EuroParl koehn2005europarl 7,457 7.97 176.14 70.24
Fisher cieri2004fisher 19,335 37.17 240.00 135.87
Switchboard godfrey1992switchboard 1,352 22.81 148.96 87.40
MultiDialog mesaros2018multi 20,811 3.17 499.33 135.62
DailyTalk lee2023dailytalk 7,368 8.05 103.66 31.15
VoxCeleb2 chung2018voxceleb2 6,171 8.06 1193.60 71.08
@@ Appendix F Prompts
We provide all prompting templates used across our datasets and QA types in Figures 14, 15, 16, 17, 18, 19, 20, 21, 22, 23, 24, 25, 26, 27, 28, 29, 30, 31, 32, 33, 34, 35, 36, 37, 38, 39, 40, 41, 42, 43, 44, 45, 46, 47, 48, and 49.
@@ Appendix G AF3 Training Datasets
Table 11 summarizes all datasets used to train AF3, including total hours, number of audio-QA pairs, and the number of epochs (passes over the dataset) used at each training stage. Similar to  (kong2025audioflamingo2, ), we convert all foundational datasets (captioning, classification, etc.) into QA formats, using the same set of prompts for each task mentioned in  (kong2025audioflamingo2, ).
@@
LongAudioXL (Ours) - 1000K $1.0$ $1.0$ $1.0$ $1.0$
AF-Think (Ours) - 250K $1.0$ $1.0$ $1.0$ $2.0$
AF-Chat (Ours) - 75K - - - - $1.0$
CompA-R ghoshcompa 159 hrs 350k - 2.0 2.0 - -
MusicBench melechovsky2023mustango 115.5 hrs 686k - 1.0 1.0 - -
Mu-LLAMA liu2024music 62.9 hrs 70k 1.0 2.0 2.0 - -
Salmonn AQA tang2023salmonn 800 hrs 270k - 1.0 1.0 - -
ClothoAQA lipping2022clotho 7.4 hrs 9.7K - $8.0$ $8.0$
OpenAQA gong2023ltu 693.2 hrs 1959.8K - $1.0$ $1.0$
Clotho-v2 drossos2020clotho 24.0 hrs 19.2K 1.0 $2.0$ $2.0$
MACS morato2021diversity 10.9 hrs 17.3K - $1.0$ $1.0$
FSD50k fonseca2021fsd50k 80.8 hrs 41.0K 1.0 $1.0$ $1.0$
CochlScene jeong2022cochlscene 169.0 hrs 60.9K - $1.0$ $1.0$
NonSpeech 7k rashid2023nonspeech7k 6.2 hrs 6.3K - $4.0$ $4.0$
Chime-home foster2015chime 5.0 hrs 4.5K - $1.0$ $1.0$
Sonyc-UST cartwright2020sonyc 34.9 hrs 27.9K - $1.0$ $1.0$
Emov-DB moon2022language 7.8 hrs 6.8K - $1.0$ $1.0$
JL-Corpus james2018open 1.4 hrs 2.4K - $6.0$ $6.0$
Tess 1.6 hrs 2.8K - $2.0$ $2.0$
OMGEmotion barros2018omg 3.0 hrs 1.7K - $3.0$ $3.0$
MusicAVQAaudio-only li2022learning 77.1 hrs 5.7K - $6.0$ $6.0$
MusicQA ouyang2025mqad 62.9 hrs 70K - $1.0$ $1.0$
LP-MusicCapsMSD doh2023lp 5805.7 hrs 1331.8K 1.0 $1.0$ $1.0$
LP-MusicCapsMTT doh2023lp 126.4 hrs 46.9K 1.0 $1.0$ $1.0$
LP-MusicCapsMC doh2023lp 7.4 hrs 7.9K 1.0 $2.0$ $2.0$
MusicCaps agostinelli2023musiclm 7.4 hrs 2.6K 1.0 $6.0$ $6.0$
NSynth engel2017neural 321.3 hrs 289.2K - $8.0$ $8.0$
MusDB-HQ rafii2017musdb18 29.1 hrs 10.2K - $2.0$ $2.0$
FMA defferrard2016fma 860.7 hrs 104.2K - $1.0$ $1.0$
Laion630kBBCSoundEffects wu2023large 456.9 hrs 15.1K $1.0$ $1.0$
Laion630kFreesound wu2023large 2494.8 hrs 306.5K $1.0$ $1.0$
SoundDescs koepke2022audio 749.7 hrs 23.1K $1.0$ $1.0$
WavCaps mei2024wavcaps 3793.3 hrs 402.6 K $1.0$ $1.0$
AudioSet gemmeke2017audio 2617.8 hrs 950.8K 1.0 - $1.0$
WavText5K deshmukh2022audio 23.8 hrs 4.3K $1.0$ $1.0$
MSP-Podcast martinez2020msp 73.9 hrs 45.1K $1.0$ $1.0$ $1.0$
MELD poria2018meld 8.7 hrs 32.9K $1.0$ $1.0$ $1.0$
MusicAVQAaudio-visual li2022learning 142.4 hrs 17.9K $1.0$ $6.0$ $6.0$
Music4All Captions (ours) 910.5 hrs 55.6K $1.0$ $1.0$
MSD Captions (ours) 15449.9 hrs 55.6K $1.0$ $1.0$
Speech-in-Sound Captions (ours) 6227.6 hrs 1999959 $1.0$ $1.0$
LibriSpeech panayotov2015librispeech 960 hrs 281.2K $1.0$ $1.0$ $1.0$
Switchboard godfrey1992switchboard 109.9 hrs 76.6K $1.0$ $1.0$ $1.0$
GigaSpeech (L) chen2021gigaspeech 2499.8 hrs 2266.3K $1.0$ $1.0$ $1.0$
Common Voice 15 commonvoice:2020 1752.1 hrs 1109.6K $1.0$ $1.0$ $1.0$
VoxPopuli (en) wang2021voxpopuli 501.8 hrs 177K $1.0$ $1.0$ $1.0$
TEDLIUM (en) hernandez2018ted 472.3 hrs 68K $1.0$ $1.0$ $1.0$
SPGISpeech o2021spgispeech 4999.8 hrs 1966.1K $1.0$ $1.0$ $1.0$
VoiceAssistant400K (xie2024mini, ) 684 hrs 470K - - - - $1.0$

### v2

@@ A.2 Training Datasets
GigaSpeech (L) [14] 2,266,371
Speech-in-Sound Captions [2] 1,999,959
SPGISpeech [90] 1,966,109
Sound-VECaps [119] 1,657,029
Million Songs Dataset [9] 1,169,997
Common Voice 15 [5] 1,109,689
MiraData [58] 748,320
Action2sound [13] 306,602
NSynth [34] 289,205
LibriSpeech [92] 281,241
Freesound [37] 256,695
AudioSet Strong [50] 216,622
VGGSound [15] 185,161
VoxPopuli (en) [107] 177,019
FMA [25] 106,412
Video Recap [55] 64,627
CochlScene [57] 60,855
Music4All [102] 109269
Switchboard [43] 76,652
FSD50k [36] 40,966
MACS [87] 31,675
@@
MagnaTagATune [67] 25,863
SoundDescs [63] 23,085
Clotho [32] 19,195
TAU-Urban [85] 14,400
MusicCaps [3] 5,479
WavText5K [28] 4,347
SONICS [98] 1,602
@@
MUSDB18 [97] 276
Medleydb-Pitch [10] 103
@@ Appendix B AudioSkills-XL
Temporal 188K Table 14 in [40] Table 14 in [40] Table 14 in [40]
@@
Attribute Identification 201K Table 14 in [40] Table 14 in [40] Table 14 in [40]
Counting 50K Table 14 in [40] Table 14 in [40] Table 14 in [40]
Contextual Sound Event Reasoning 982K Table 14 in [40] Table 14 in [40] Table 14 in [40]
Contextual Speech Event Reasoning 1,272K Table 14 in [40] Table 14 in [40] Table 14 in [40]
Information Extraction 858K Table 14 in [40] Table 14 in [40] Table 14 in [40]
General Reasoning 704K Table 14 in [40] Table 14 in [40] Table 14 in [40]
@@
Speech-in-Sound QA (ours) 1,739K Speech-in-Sound Caps (YouTube8M) Caption, Transcripts, Speech Characteristics Fig. 26 50
@@ E.2 Clustering for constructing AF-Chat
Specifically, for each seed audio, we retrieve its top 8 most semantically similar and top 8 most dissimilar clips from the dataset. Similarity is computed using captions, NV-Embed-v2 embeddings of the captions, and FAISS-based similarity search [31] of the embeddings.
@@
Order VoxPopuli [107] 16,926 1.87 294.80 89.55
LibriSpeech [92] 2,340 16.02 147.59 82.01
MELD [94] 4,135 1.06 108.01 30.42
IEMOCAP [11] 599 82.00 542.00 272.45
EuroParl [62] 11,885 2.59 176.14 69.34
Fisher [22] 25,962 33.34 240.00 136.84
Switchboard [43] 2,702 22.81 148.96 87.38
MultiDialog [85] 27,927 1.31 499.33 135.10
VoxCeleb2 [21] 12,855 8.00 1273.60 71.12
Emotion Ident. IEMOCAP [11] 300 82.00 542.00 272.22
MELD [94] 1,847 1.78 108.01 33.20
Emotion Causal Reason. IEMOCAP [11] 300 82.00 542.00 272.22
MELD [94] 1,850 1.57 108.01 33.13
Emotion Flip Reason. IEMOCAP [11] 299 82.00 542.00 272.62
MELD [94] 1,807 1.64 108.01 33.57
Topic Relation. Reason. VoxPopuli [107] 13,651 3.58 240.44 97.14
LibriSpeech [92] 1,165 16.02 147.59 82.11
MELD [94] 1,518 1.89 108.01 34.45
IEMOCAP [11] 188 82.00 542.00 270.99
EuroParl [62] 9,381 7.97 176.14 70.14
Fisher [22] 20,453 33.34 240.00 136.10
Switchboard [43] 998 24.58 148.96 90.05
MultiDialog [85] 14,906 5.11 499.33 135.35
DailyTalk [71] 3,141 8.05 103.66 35.32
VoxCeleb2 [21] 5,414 8.51 1193.60 78.96
Sarcasm Ident. IEMOCAP [11] 299 82.00 542.00 271.58
MELD [94] 1,958 1.10 108.01 31.82
Summarization VoxPopuli [107] 13,913 2.12 294.80 91.38
LibriSpeech [92] 1,057 16.02 147.59 83.15
MELD [94] 2,803 1.84 108.01 32.92
IEMOCAP [11] 300 82.00 542.00 272.22
EuroParl [62] 8,905 6.62 176.14 70.03
Fisher [22] 15,500 0.33 240.00 135.60
Switchboard [43] 1,346 24.58 148.96 87.60
MultiDialog [85] 20,838 1.93 499.33 135.73
DailyTalk [71] 7,218 8.05 103.66 31.42
VoxCeleb2 [21] 5,894 7.94 1193.60 70.87
Spotify Podcasts [23] 103920 0.06 18206.44 2002.99
Needle QA (IE) DailyTalk [71] 13,563 5.72 103.66 31.12
EuroParl [62] 18,426 6.57 176.14 70.10
Fisher [22] 37,779 18.59 240.00 135.99
IEMOCAP [11] 542 82.00 542.00 272.03
LibriSpeech [92] 2,248 16.02 147.59 82.82
Spotify Podcasts [23] 103920 0.06 18206.44 2002.99
Response QA (IE) VoxPopuli [107] 13,913 2.12 294.80 91.38
MELD [94] 1,660 1.57 108.01 31.83
IEMOCAP [11] 177 82.00 542.00 272.52
MultiDialog [85] 13,505 1.95 499.33 135.40
DailyTalk [71] 4,516 5.72 103.66 30.91
Switchboard [43] 862 22.81 148.96 88.75
@@
Causal QA (IE) VoxPopuli [107] 12,264 4.10 240.44 92.88
LibriSpeech [92] 1,166 16.02 147.59 82.04
MELD [94] 2,957 1.27 108.01 31.74
IEMOCAP [11] 298 82.00 542.00 273.10
EuroParl [62] 7,457 7.97 176.14 70.24
Fisher [22] 19,335 37.17 240.00 135.87
Switchboard [43] 1,352 22.81 148.96 87.40
MultiDialog [85] 20,811 3.17 499.33 135.62
DailyTalk [71] 7,368 8.05 103.66 31.15
VoxCeleb2 [21] 6,171 8.06 1193.60 71.08
@@ Appendix F Prompts
We provide all prompting templates used across our datasets and QA types in Figures 15, 16, 17, 18, 19, 20, 21, 22, 23, 24, 25, 26, 27, 28, 29, 30, 31, 32, 33, 34, 35, 36, 37, 38, 39, 40, 41, 42, 43, 44, 45, 46, 47, 48, 49, and 50.
@@ Appendix G AF3 Training Datasets
Table 11 summarizes all datasets used to train AF3, including total hours, number of audio-QA pairs, and the number of epochs (passes over the dataset) used at each training stage. Similar to  [40], we convert all foundational datasets (captioning, classification, etc.) into QA formats, using the same set of prompts for each task mentioned in  [40].
@@
LongAudioXL (Ours) - 1000K 1.01.01.0 1.01.01.0 1.01.01.0 1.01.01.0 -
AF-Think (Ours) - 250K 1.01.01.0 1.01.01.0 1.01.01.0 2.02.02.0 -
AF-Chat (Ours) - 75K - - - - 1.01.01.0
CompA-R [42] 159 hrs 350k - 2.0 2.0 - -
MusicBench [84] 115.5 hrs 686k - 1.0 1.0 - -
Mu-LLAMA [78] 62.9 hrs 70k 1.0 2.0 2.0 - -
Salmonn AQA [104] 800 hrs 270k - 1.0 1.0 - -
ClothoAQA [76] 7.4 hrs 9.7K - 8.08.08.0 8.08.08.0 - -
OpenAQA [46] 693.2 hrs 1959.8K - 1.01.01.0 1.01.01.0 - -
Clotho-v2 [32] 24.0 hrs 19.2K 1.0 2.02.02.0 2.02.02.0 - -
MACS [87] 10.9 hrs 17.3K - 1.01.01.0 1.01.01.0 - -
FSD50k [36] 80.8 hrs 41.0K 1.0 1.01.01.0 1.01.01.0 - -
CochlScene [57] 169.0 hrs 60.9K - 1.01.01.0 1.01.01.0 - -
NonSpeech 7k [99] 6.2 hrs 6.3K - 4.04.04.0 4.04.04.0 - -
Chime-home [38] 5.0 hrs 4.5K - 1.01.01.0 1.01.01.0 - -
Sonyc-UST [12] 34.9 hrs 27.9K - 1.01.01.0 1.01.01.0 - -
Emov-DB [86] 7.8 hrs 6.8K - 1.01.01.0 1.01.01.0 - -
JL-Corpus [56] 1.4 hrs 2.4K - 6.06.06.0 6.06.06.0 - -
Tess 1.6 hrs 2.8K - 2.02.02.0 2.02.02.0 -
OMGEmotion [7] 3.0 hrs 1.7K - 3.03.03.0 3.03.03.0 - -
MusicAVQAaudio-only [74] 77.1 hrs 5.7K - 6.06.06.0 6.06.06.0 - -
MusicQA [91] 62.9 hrs 70K - 1.01.01.0 1.01.01.0 - -
LP-MusicCapsMSD [30] 5805.7 hrs 1331.8K 1.0 1.01.01.0 1.01.01.0 - -
LP-MusicCapsMTT [30] 126.4 hrs 46.9K 1.0 1.01.01.0 1.01.01.0 - -
LP-MusicCapsMC [30] 7.4 hrs 7.9K 1.0 2.02.02.0 2.02.02.0 - -
MusicCaps [3] 7.4 hrs 2.6K 1.0 6.06.06.0 6.06.06.0 - -
NSynth [34] 321.3 hrs 289.2K - 8.08.08.0 8.08.08.0 - -
MusDB-HQ [97] 29.1 hrs 10.2K - 2.02.02.0 2.02.02.0 - -
FMA [25] 860.7 hrs 104.2K - 1.01.01.0 1.01.01.0 - -
Laion630kBBCSoundEffects [113] 456.9 hrs 15.1K 1.01.01.0 - 1.01.01.0 - -
Laion630kFreesound [113] 2494.8 hrs 306.5K 1.01.01.0 - 1.01.01.0 - -
SoundDescs [63] 749.7 hrs 23.1K 1.01.01.0 - 1.01.01.0 - -
WavCaps [83] 3793.3 hrs 402.6 K 1.01.01.0 - 1.01.01.0 - -
AudioSet [39] 2617.8 hrs 950.8K 1.0 - 1.01.01.0 - -
WavText5K [28] 23.8 hrs 4.3K 1.01.01.0 - 1.01.01.0 - -
MSP-Podcast [82] 73.9 hrs 45.1K 1.01.01.0 1.01.01.0 1.01.01.0 - -
MELD [94] 8.7 hrs 32.9K 1.01.01.0 1.01.01.0 1.01.01.0 - -
MusicAVQAaudio-visual [74] 142.4 hrs 17.9K 1.01.01.0 6.06.06.0 6.06.06.0 - -
Music4All Captions (ours) 910.5 hrs 55.6K 1.01.01.0 - 1.01.01.0 - -
MSD Captions (ours) 15449.9 hrs 55.6K 1.01.01.0 - 1.01.01.0 - -
Speech-in-Sound Captions (ours) 6227.6 hrs 1999959 1.01.01.0 - 1.01.01.0 - -
LibriSpeech [92] 960 hrs 281.2K 1.01.01.0 1.01.01.0 1.01.01.0 - -
Switchboard [43] 109.9 hrs 76.6K 1.01.01.0 1.01.01.0 1.01.01.0 - -
GigaSpeech (L) [14] 2499.8 hrs 2266.3K 1.01.01.0 1.01.01.0 1.01.01.0 - -
Common Voice 15 [5] 1752.1 hrs 1109.6K 1.01.01.0 1.01.01.0 1.01.01.0 - -
VoxPopuli (en) [107] 501.8 hrs 177K 1.01.01.0 1.01.01.0 1.01.01.0 - -
TEDLIUM (en) [49] 472.3 hrs 68K 1.01.01.0 1.01.01.0 1.01.01.0 - -
SPGISpeech [90] 4999.8 hrs 1966.1K 1.01.01.0 1.01.01.0 1.01.01.0 - -
VoiceAssistant400K [115] 684 hrs 470K - - - - 1.01.01.0